# Variations on quantum de Finetti theorems and operator valued Martin boundaries: a Choquet–Deny approach

### v2

@@ Abstract.
We revisit the quantum de Finetti theorem. We state and prove a couple of variants thereof. In parallel, we introduce an operator version of the Martin boundary on quantum groups and prove generalizations of Biane’s theorem. Our proof of the de Finetti theorem is new in the sense that it is based on an analogy with the theory of operator valued Martin boundary that we introduce.
@@ 1. Introduction
The quantum de Finetti theorem is a fundamental theorem in quantum information theory. It states that the collection of separable states corresponds precisely to the collection of infinitely extendable symmetric states. It was developed independently in operator algebras [12] and in quantum information theory [7]. This remains one of the most powerful tools to detect entanglement: a state is entangled if and only if it fails to be $k$-extendable for $k$ large enough. In general, proofs of this phenomenon rely heavily on the representation theory of the symmetric group.
In this paper, we propose a new approach to this theorem, inspired by the theory of random walks on commutative groups, and, more precisely, on the notion of the Martin boundary and the Choquet-Deny theory. In these theories, the harmonic elements, which are considered, are scalar-valued functions because studying more general functions do not enrich the theory. In their non-commutative generalizations, Biane and Izumi did not consider operator-valued boundaries. In this paper, we study operator-valued boundaries and observe that it is closely related to the problem of identifying separable states.
After a short section on notations, we provide in Section 3 a new proof of the generalized quantum de Finetti theorem, inspired by harmonic analysis. It mimics Biane’s proof of the non-commutative Choquet-Deny theorem.
@@
In Section 5, we explore further the relation between sections 3 and 4 by showing an inclusion between the matrix-valued non-commutative Martin boundary and the collection of extendable matrix-valued states. This inclusion is based on a matrix-valued extension of a non-commutative Martingale convergence theorem considered by Biane and Izumi.
The last section delves into the problem of understanding the notion of exponentials better. We extend some results of Biane [2] to more groups, and in the matrix-valued setup. This allows us to identify cases when the injection described in part 3 is a bijection.
@@ 2. Notations
In this section, we summarize the notations and conventions that we use throughout this paper. Let $\mathbb{M}_{n}(\mathbb{C})$ denote the *-algebra of $n\times n$ matrices over $\mathbb{C}$ and $\mathrm{Tr}_{n}$ the (non-normalized) trace on $\mathbb{M}_{n}(\mathbb{C})$. We denote by $\mathbb{M}_{n}(\mathbb{C})_{+}$ the set of all positive semi-definite matrices. In the paper, elements in $\mathbb{M}_{n}(\mathbb{C})_{+}$ are simply called positive. More generally, let $B(\mathcal{H})$ denote the *-algebra of bounded linear operators on a complex Hilbert space $\mathcal{H}$. Here, its multiplication is given as compositions of linear operators and its *-operation is given by taking adjoint operators. Let $\mathcal{A}\subseteq B(\mathcal{H})$ be a unital $C^{*}$-algebra, i.e., $\mathcal{A}$ is a unital *-subalgebra of $B(\mathcal{H})$ and closed with respect to the topology induced by the operator norm. We denote by $\mathcal{A}_{+}$ the set of all positive elements in $\mathcal{A}$, that is, by definition, $x\in\mathcal{A}_{+}$ if and only if $x$ is self-adjoint and the spectrum of $x$ is contained in $\mathbb{R}_{\geq 0}$. For any pair $x,y\in\mathcal{A}_{+}$ we write $x\leq y$ if $y-x\in\mathcal{A}_{+}$. We remark that $\mathcal{A}_{+}$ is a convex cone.
Let $\mathcal{A}^{*}$ denote the set of all (continuous) linear functionals on $\mathcal{A}$. We say that $\rho\in\mathcal{A}^{*}$ is positive if $\rho(x)\geq 0$ for any $x\in\mathcal{A}_{+}$, and $\mathcal{A}^{*}_{+}$ denotes the set of all positive linear functionals on $\mathcal{A}$. In addition, if $\rho(1_{\mathcal{A}})=1$ holds, then $\rho$ is said to be a state of $\mathcal{A}$. Here, $1_{\mathcal{A}}$ is the multiplicative unit in $\mathcal{A}$. We say that $\rho\in\mathcal{A}^{*}_{+}$ is faithful if $\rho(x)>0$ for any $x\in\mathcal{A}_{+}\backslash\{0\}$.
@@
If $\mathcal{H}$ is finite-dimensional, then $B(\mathcal{H})$ coincides with $\mathrm{End}(\mathcal{H})$ (and $\mathbb{M}_{n}(\mathbb{C})$ if $\dim\mathcal{H}=n$). Namely, all linear operator on $\mathcal{H}$ are bounded. Moreover, the trace $\mathrm{Tr}$ on $B(\mathcal{H})$ can be defined. Using $\mathrm{Tr}$, we define the Hilbert–Schmidt inner product on $B(\mathcal{H})$ by $\langle A,B\rangle:=\mathrm{Tr}(B^{*}A)$ for all $A,B\in B(\mathcal{H})$. Moreover, by the Hilbert–Schmidt inner product, $B(\mathcal{H})$ can be naturally identified with its dual space $B(\mathcal{H})^{*}$.
We also use the following terminologies in convex analysis (see e.g., [11]). Let $K$ be a convex cone and $\leq$ denote the induced partial order on the vector space $E:=K-K$. A subset $\alpha\subset K$ is called a ray in $K$ if $\alpha=\{rx\mid r\geq 0\}$ for some $x\in K$. A ray $\alpha$ is said to be extreme if $ry\leq x$ implies $y\in\alpha$ for any $x\in\alpha$, $y\in K$, and $r\geq 0$. An element of extreme ray is said to be extreme in $K$, and $\mathrm{ex}(K)$ denotes the set of all extreme elements in $K$. By definition, $x\in\mathrm{ex}(K)$ holds if and only if $y\leq x$ implies $y=rx$ for some $r\geq 0$ for any $y\in K$. We suppose that $E$ is locally convex. We say that a nonnegative Borel measure $M$ on $K$ represents a point $x\in K$ such that $f(x)=\int_{K}f(y)M(dy)$ for any $f\in E^{*}$. For any two nonnegative Borel measures $M_{1},M_{2}$ on $K$ we write $M_{1}\prec M_{2}$ if $\int_{K}f(x)M_{1}(dx)\leq\int_{K}f(x)M_{2}(x)$ for any convex function $f$ on $K$.
We assume that $K$ is closed. A non-empty subset $C$ of $K$ is called a cap of $K$ if $C$ is compact, convex, and $K\backslash C$ is convex. See [11, p.80].
@@ 3. The generalized quantum de Finetti theorem via Choquet theory
Throughout this section, we fix two natural numbers $n,m$ and $\rho\in\mathbb{M}_{n}(\mathbb{C})^{*}_{+}$ that is faithful. The identity map on $\mathbb{M}_{n}(\mathbb{C})$ is simply denoted by $\mathrm{id}_{n}$. We define
@@
where $\mathbb{M}_{n}(\mathbb{C})^{\otimes l}$ is the tensor product of $l$ copies of $\mathbb{M}_{n}(\mathbb{C})$ for $l\geq 1$, and $\mathbb{M}_{n}(\mathbb{C})^{\otimes 0}:=\mathbb{C}$. Moreover, $S_{l}$ is the symmetric group of degree $l$, and it acts on $\mathbb{M}_{n}(\mathbb{C})^{\otimes l}$ by
We denote by $(\mathbb{M}_{n}(\mathbb{C})^{\otimes l})^{S_{l}}$ the *-subalgebra of fixed points under this action. We equip $K_{m,\rho}$ with the topology of component-wise weak convergence, where $(\mathbb{M}_{n}(\mathbb{C})^{\otimes l})^{S_{l}}$ is identified with the subspace of the $S_{l}$-invariant elements in $(\mathbb{M}_{n}(\mathbb{C})^{\otimes l})^{*}$ by the Hilbert–Schmidt inner product. We also remark that $K_{m,\rho}$ is a convex cone by component-wise operations.
@@
$K_{m,\rho}$ is the closed convex hull of $\mathrm{ex}(K_{m,\rho})$.
@@
For any $l\geq 1$ and $x\in C_{r}$ we have $(\mathrm{Tr}_{m}\otimes\rho^{\otimes l})(|x_{l}|)=(\mathrm{Tr}_{m}\otimes\rho^{\otimes l})(x_{l})\leq\mathrm{Tr}_{m}(x_{0})\leq 1$. Thus, $C_{r}$ is compact, and hence $C_{r}$ is a cap of $K_{m,\rho}$ (see [11, Proposition 13.2]). Moreover, $K_{m,\rho}=\bigcup_{r>0}C_{r}$ holds. Thus, the assertion follows from [11, Choquet theorem in p.81]. ∎
We remark that the above proof showed that $K_{m,\rho}$ is a union of its caps. Thus, combining with results in [11, Chapter 13], the following representation theorem holds true:
@@
Every $x\in K_{m,\rho}\backslash\{0\}$ can be represented by a Borel probability measure $M_{x}$ on $K_{m,\rho}$ such that $M_{x}$ is maximal with respect to the partial order $\prec$ and supported on $\mathrm{ex}(K_{m,\rho})$ (i.e., $M_{x}(K_{m,\rho}\backslash\mathrm{ex}(K_{m,\rho}))=0$).
We obtain the following description of extreme points of $K_{m,\rho}$.
@@
$\mathrm{ex}(K_{m,\rho})$ is contained in $\{(a\otimes b^{\otimes l})_{l=0}^{\infty}\mid a\in\mathbb{M}_{m}(\mathbb{C})_{+},b\in\mathbb{M}_{n}(\mathbb{C})_{+},\rho(b)\leq 1\}$.
@@
Let us fix $x=(x_{l})_{l=0}^{\infty}\in\mathrm{ex}(K_{m,\rho})$. For any $\nu\in\mathbb{M}_{n}(\mathbb{C})^{*}_{+}$ we define
@@
Thus, $x_{\nu}\leq x$ holds, and hence there exists $b(\nu)\geq 0$ such that $x_{\nu}=b(\nu)x$ since $x\in\mathrm{ex}(K_{m,\rho})$. We claim that the mapping from $\nu$ to $b(\nu)$ can be extended linearly to a functional on $\mathbb{M}_{n}(\mathbb{C})^{*}$. Let $\nu\in\mathbb{M}_{n}(\mathbb{C})^{*}_{+}$ (not necessarily $\nu\leq\rho$). Since $\rho$ is faithful on $\mathbb{M}_{n}(\mathbb{C})$, there exists $r>0$ such that $r\nu\leq\rho$. Thus, for some $b(r\nu)>0$, we have $x_{r\nu}=b(r\nu)x$ and $rx_{\nu}=x_{r\nu}=b(r\nu)x$. Namely, $b(r\nu)/r$ depends only on $\nu$, and hence $b(\nu):=rb(\nu/r)$ is well defined. Moreover, the mapping $b\colon\mathbb{M}_{n}(\mathbb{C})^{*}_{+}\to\mathbb{R}_{\geq 0}$ is additive and positive-homogeneous. By the Jordan decomposition (see e.g., [10]), every linear functional on $\mathbb{M}_{n}(\mathbb{C})$ uniquely decomposes into a linear combination of four positive linear functionals. Thus, the mapping $b$ can be extended to a linear functional on $\mathbb{M}_{n}(\mathbb{C})$, and $b$ is naturally in $\mathbb{M}_{n}(\mathbb{C})=(\mathbb{M}_{n}(\mathbb{C})^{*})^{*}$. Moreover, we have $b\in\mathbb{M}_{n}(\mathbb{C})_{+}$, and $(\mathrm{id}_{m}\otimes\mathrm{id}_{n}^{\otimes l}\otimes\nu)(x_{l+1})=(\mathrm{id}_{m}\otimes\mathrm{id}_{n}^{\otimes l}\otimes\nu)(x_{l}\otimes b)$ for any $l\geq 0$ and $\nu\in\mathbb{M}_{n}(\mathbb{C})^{*}$. It implies that $x_{l+1}=x_{l}\otimes b$, and hence, iterating over $l$, we have $x_{l}=x_{0}\otimes b^{\otimes l}$. ∎
Let us recall that $a\in(\mathbb{M}_{m}(\mathbb{C})\otimes\mathbb{M}_{n}(\mathbb{C}))_{+}$ is called separable if $a$ is in the convex hull of $\mathbb{M}_{m}(\mathbb{C})_{+}\otimes\mathbb{M}_{n}(\mathbb{C})_{+}$. Based on the results in this section, we obtain a characterization of separable elements, which is a generalization of the quantum de Finetti theorem.
@@
Let $\rho\in\mathbb{M}_{n}(\mathbb{C})^{*}_{+}$ and $a\in(\mathbb{M}_{m}(\mathbb{C})\otimes\mathbb{M}_{n}(\mathbb{C}))_{+}$. For every $l\geq 1$ we call $b$ in $(\mathbb{M}_{m}(\mathbb{C})\otimes(\mathbb{M}_{n}(\mathbb{C})^{\otimes l})^{S_{l}})_{+}$ a $l$-sub-extension of $a$ (with respect to $\rho$) if $(\mathrm{id}_{m}\otimes\mathrm{id}_{n}\otimes\rho^{\otimes(l-1)})(b)\leq a$ holds. If $a$ has a $l$-sub-extension, then $a$ is said to be $l$-sub-extendable (with respect to $\rho$).
@@
    $a$ is $l$-sub-extendable with respect to every faithful linear functional on $\mathbb{M}_{n}(\mathbb{C})$ and $l\geq 1$.
@@
Let $\rho,\nu\in\mathbb{M}_{n}(\mathbb{C})_{+}^{*}$ be faithful. Then, there exists $r>0$ such that $r\nu\leq\rho$. Thus, if $b$ is a $l$-sub-extension of $a$ with respect to $\rho$, then $r^{l-1}b$ is a $l$-sub-extension of $a$ with respect to $\nu$. Namely, (2), (3) are equivalent.
We show that (1), (2) are equivalent. If $a$ is separable, then it is clearly $l$-sub-extendable for all $l\geq 1$. Conversely, we assume that there exist $l$-sub-extensions $x_{l}$ of $a$ for all $l\geq 1$. For any $k\geq l$ we define

### v3

@@ Abstract.
We revisit the quantum de Finetti theorem. We state and prove a couple of variants thereof. In parallel, we introduce an operator version of the Martin boundary on quantum groups and prove generalizations of Biane’s theorem. Our proof of the de Finetti theorem is new in the sense that it is based on an analogy with the theory of operator-valued Martin boundary that we introduce.
@@ 1. Introduction
The quantum de Finetti theorem is a fundamental theorem in quantum information theory. It states that the collection of separable states corresponds precisely to the collection of infinitely symmetrically extendable states. It was developed independently in operator algebras [12] and in quantum information theory [7]. This remains one of the most powerful tools for detecting bipartite entanglement.
The classical de Finetti theorem is a cornerstone of probability theory, asserting that an infinitely exchangeable sequence of random variables is a convex mixture of independent and identically distributed (i.i.d.) random variables. This profound result has found its analogues and generalizations in various mathematical fields. Parallel to this, the theory of random walks on classical groups features the concept of a Martin boundary, which provides a geometric and probabilistic framework for characterizing positive harmonic functions—functions invariant under the transition operator of a random walk—and positive sub-harmonic functions. The Choquet–Deny theorem establishes that such harmonic functions can be uniquely represented as integrals over this boundary. Traditionally, these classical theories primarily concern scalar-valued functions. We refer the reader to [13] as a general reference on this subject.
In the quantum realm, the quantum de Finetti theorem extends this notion to quantum states, asserting that quantum states of a bipartite quantum system that are symmetrically extendable to arbitrarily larger quantum systems can be approximated by convex combinations of tensor product states. Namely, such states are separable. This finite version of the theorem is particularly vital in quantum information theory, serving as a powerful tool for entanglement detection: a state is entangled if it fails to be $l$-extendable for sufficiently large $l$. Here, $l$ presents the size of extended quantum systems. For the definition and a bit of context, we refer to Definition 3.1.
In this paper, we propose a new approach to (quantum) de Finetti-type theorems, inspired by the theory of random walks on commutative groups, and, more precisely, on the notion of the Martin boundary and the Choquet-Deny theory.
While conventional proofs of quantum de Finetti theorems often rely on the representation theory of symmetric and unitary groups, our work offers a novel approach by drawing inspiration from non-commutative harmonic analysis and the Martin boundary theory for quantum random walks, pioneered by Biane [1, 2, 3] and Izumi [6].
In these theories, the harmonic elements, which are considered, are scalar-valued functions because studying more general functions do not enrich the theory. In their non-commutative generalizations, Biane in [3] and Izumi in [6] did not consider operator-valued boundaries. In this paper, we study operator-valued boundaries and observe that it is closely related to the problem of identifying separable states.
Specifically, we introduce matrix-valued Martin boundaries and sub-harmonic elements, which generalize their scalar-valued counterparts and are shown to be intimately linked to the concept of separable quantum states of a bipartite quantum system. Our notion of $k$-sub-extendability, central to our generalized quantum de Finetti theorem (Theorem 3.2), is directly inspired by these non-commutative sub-harmonic elements. This paper thus establishes a new connection between quantum de Finetti theorems and the operator-valued Martin boundary theory.
After a short section on notations, we provide in Section 3 a new proof of the generalized quantum de Finetti theorem, inspired by harmonic analysis. It mimics Biane’s proof of the non-commutative Choquet-Deny theorem. See Theorem 3.2.
@@
In Section 5, we explore further the relation between sections 3 and 4 by showing an inclusion between the matrix-valued non-commutative Martin boundary and the collection of extendable matrix-valued states (see Proposition 5.1). This inclusion is based on a matrix-valued extension of a non-commutative Martingale convergence theorem considered by Biane and Izumi.
The last section delves into the problem of understanding the notion of exponentials better. We extend some results of Biane [2] to more groups, and in the matrix-valued setup. See Lemma 6.1. This allows us to identify cases when the injection described in part 3 is a bijection.
@@ 2. Notations
In this section, we summarize the notations and conventions used throughout this paper. Let $\mathbb{M}_{n}(\mathbb{C})$ denote the *-algebra of $n\times n$ matrices over $\mathbb{C}$ and $\mathrm{Tr}_{n}$ the (non-normalized) trace on $\mathbb{M}_{n}(\mathbb{C})$. We denote by $\mathbb{M}_{n}(\mathbb{C})_{+}$ the set of all positive semi-definite matrices. In the paper, elements in $\mathbb{M}_{n}(\mathbb{C})_{+}$ are simply called positive. More generally, let $B(\mathcal{H})$ denote the *-algebra of bounded linear operators on a complex Hilbert space $\mathcal{H}$. Here, its multiplication is given as compositions of linear operators and its *-operation is given by taking adjoint operators. Let $\mathcal{A}\subseteq B(\mathcal{H})$ be a unital $C^{*}$-algebra, i.e., $\mathcal{A}$ is a unital *-subalgebra of $B(\mathcal{H})$ and closed with respect to the topology induced by the operator norm. We denote by $\mathcal{A}_{+}$ the set of all positive elements in $\mathcal{A}$, that is, by definition, $x\in\mathcal{A}_{+}$ if and only if $x$ is self-adjoint and the spectrum of $x$ is contained in $\mathbb{R}_{\geq 0}$. For any pair $x,y\in\mathcal{A}_{+}$ we write $x\leq y$ if $y-x\in\mathcal{A}_{+}$. We remark that $\mathcal{A}_{+}$ is a proper convex cone and that as $\mathcal{A}_{+}-\mathcal{A}_{+}=\mathcal{A}_{\mathrm{sa}}$ and $\mathcal{A}_{+}\cap(-\mathcal{A}_{+})=\{0\}$, the self-adjoint elements $\mathcal{A}_{\mathrm{sa}}$ of $\mathcal{A}$ becomes a partially ordered real vector space by defining $x\leq y$ if $y-x\in\mathcal{A}_{+}$.
[…]
@@
If $\mathcal{H}$ is finite-dimensional, then $B(\mathcal{H})$ coincides with $\mathrm{End}(\mathcal{H})$ (and $\mathbb{M}_{n}(\mathbb{C})$ if $\dim\mathcal{H}=n$). Namely, all linear operators on $\mathcal{H}$ are bounded. Moreover, the trace $\mathrm{Tr}$ on $B(\mathcal{H})$ can be defined. Using $\mathrm{Tr}$, we define the Hilbert–Schmidt inner product on $B(\mathcal{H})$ by $\langle A,B\rangle:=\mathrm{Tr}(B^{*}A)$ for all $A,B\in B(\mathcal{H})$. Moreover, by the Hilbert–Schmidt inner product, $B(\mathcal{H})$ can be naturally identified with its dual space $B(\mathcal{H})^{*}$.
We will use the terminology and results from convex analysis in [11], more precisely from Chapter 13 of [11]. Let $K$ be a proper convex cone (i.e., $K\cap(-K)=\{0\}$) and E:=KKE:=K-K be the associated partially ordered real vector space. Recall that a ray $\alpha$ of $K$ is a set of the form $\mathbb{R}x=\{rx\mid r\geq 0\}$, where $x\in K\backslash\{0\}$. A ray $\alpha$ is said to be an extreme ray of $K$ if whenever $x\in\alpha$ and $x=\lambda y+(1-\lambda)z$ ($y,z\in K$, $0<\lambda<1$), then $y,z\in\alpha$. We denote by $\mathrm{exr}(K)$ the union of extreme rays of $K$. Then an element $x\in K$ is in $\mathrm{exr}(K)$ if and only if $y=rx$ for some $r\geq 0$ whenever $0\leq y\leq x$.
[…]
Recall ([11, page 1]) also that the following notion which we will first use in Proposition 3.1. Let $X$ be a nonempty compact subset of a locally convex space $E$ and $\mu$ a Borel probability measure on $X$. A point $x\in E$ is represented by $\mu$ if $f(x)=\int_{X}fd\mu$ for every continuous linear functional on $E$. Moreover, if $\mu_{1}$ and $\mu_{2}$ are nonnegative regular Borel measure on $X$, then $\mu_{1}\prec\mu_{2}$ if $\int_{X}fd\mu_{1}\leq\int_{X}fd\mu_{2}$ for any convex function $f$ on $X$.
@@ 3. The generalized quantum de Finetti theorem via Choquet theory
Throughout this section, we fix two natural numbers $n,m$ and $\rho\in\mathbb{M}_{n}(\mathbb{C})^{*}_{+}$ that is faithful. The identity map on $\mathbb{M}_{n}(\mathbb{C})$ is simply denoted by $\mathrm{id}_{n}$. Let $S_{l}$ denote the symmetric group of degree $l\geq 1$, and it acts on $\mathbb{M}_{n}(\mathbb{C})^{\otimes l}$ by
We then denote by $(\mathbb{M}_{n}(\mathbb{C})^{\otimes l})^{S_{l}}$ the *-subalgebra of fixed points under this action. Moreover, the convex cone of sub-martingale sequences of symmetric quantum states is defined by
@@
where $\mathbb{M}_{n}(\mathbb{C})^{\otimes 0}:=\mathbb{C}$ and $S_{0}$ is the trivial group. We equip $K_{m,\rho}$ with the topology of component-wise weak convergence, where $(\mathbb{M}_{n}(\mathbb{C})^{\otimes l})^{S_{l}}$ is identified with the subspace of the $S_{l}$-invariant elements in $(\mathbb{M}_{n}(\mathbb{C})^{\otimes l})^{*}$ by the Hilbert–Schmidt inner product. We also remark that $K_{m,\rho}$ is a convex cone by component-wise operations.
@@
$K_{m,\rho}$ is the closed convex hull of $\mathrm{exr}(K_{m,\rho})$.
@@
For any $l\geq 1$ and $x\in C_{r}$ we have $(\mathrm{Tr}_{m}\otimes\rho^{\otimes l})(|x_{l}|)=(\mathrm{Tr}_{m}\otimes\rho^{\otimes l})(x_{l})\leq\mathrm{Tr}_{m}(x_{0})\leq r$. Thus, $C_{r}$ is compact, and hence $C_{r}$ is a cap of $K_{m,\rho}$ (see [11, Proposition 13.2]). Moreover, $K_{m,\rho}=\bigcup_{r>0}C_{r}$ holds. Thus, the assertion follows from the Choquet theorem [11, p.81]. ∎
We remark that the above proof showed that $K_{m,\rho}$ is a union of its caps. Thus, combining with results in [11, Chapter 13], we obtain the following representation theorem.
@@
Every $x\in K_{m,\rho}\backslash\{0\}$ can be represented by a Borel probability measure $M_{x}$ on $K_{m,\rho}$ such that $M_{x}$ is supported on any closed subset of $K_{m,\rho}$ which contains $\mathrm{exr}(K_{m,\rho})$.
Proof.
[…]
We obtain the following description of $\mathrm{exr}(K_{m,\rho})$.
@@
$\mathrm{exr}(K_{m,\rho})$ is contained in $\{(a\otimes b^{\otimes l})_{l=0}^{\infty}\mid a\in\mathbb{M}_{m}(\mathbb{C})_{+},b\in\mathbb{M}_{n}(\mathbb{C})_{+},\rho(b)\leq 1\}$.
@@
Let us fix $x=(x_{l})_{l=0}^{\infty}\in\mathrm{exr}(K_{m,\rho})$. For any $\nu\in\mathbb{M}_{n}(\mathbb{C})^{*}_{+}$ we define
@@
Thus, $x_{\nu}\leq x$ holds, and hence there exists $b(\nu)\geq 0$ such that $x_{\nu}=b(\nu)x$ since $x\in\mathrm{exr}(K_{m,\rho})$. We claim that the mapping from $\nu$ to $b(\nu)$ can be extended linearly to a functional on $\mathbb{M}_{n}(\mathbb{C})^{*}$. Let $\nu\in\mathbb{M}_{n}(\mathbb{C})^{*}_{+}$ (not necessarily $\nu\leq\rho$). Since $\rho$ is faithful on $\mathbb{M}_{n}(\mathbb{C})$, there exists $r>0$ such that $r\nu\leq\rho$. Thus, for some $b(r\nu)>0$, we have $x_{r\nu}=b(r\nu)x$ and $rx_{\nu}=x_{r\nu}=b(r\nu)x$. Namely, $b(r\nu)/r$ depends only on $\nu$, and hence $b(\nu):=rb(\nu/r)$ is well defined. Moreover, the mapping $b\colon\mathbb{M}_{n}(\mathbb{C})^{*}_{+}\to\mathbb{R}_{\geq 0}$ is additive and positive-homogeneous. By the Jordan decomposition (see e.g., [10]), every linear functional on $\mathbb{M}_{n}(\mathbb{C})$ uniquely decomposes into a linear combination of four positive linear functionals. Thus, the mapping $b$ can be extended to a linear functional on $\mathbb{M}_{n}(\mathbb{C})$, and $b$ is naturally in $\mathbb{M}_{n}(\mathbb{C})=(\mathbb{M}_{n}(\mathbb{C})^{*})^{*}$. Moreover, we have $b\in\mathbb{M}_{n}(\mathbb{C})_{+}$, and $(\mathrm{id}_{m}\otimes\mathrm{id}_{n}^{\otimes l}\otimes\nu)(x_{l+1})=(\mathrm{id}_{m}\otimes\mathrm{id}_{n}^{\otimes l}\otimes\nu)(x_{l}\otimes b)$ for any $l\geq 0$ and $\nu\in\mathbb{M}_{n}(\mathbb{C})^{*}$. It implies that $x_{l+1}=x_{l}\otimes b$, and hence, iterating over $l$, we have $x_{l}=x_{0}\otimes b^{\otimes l}$. ∎
Let us recall that $a\in(\mathbb{M}_{m}(\mathbb{C})\otimes\mathbb{M}_{n}(\mathbb{C}))_{+}$ is called separable if $a$ is in the convex hull of $\mathbb{M}_{m}(\mathbb{C})_{+}\otimes\mathbb{M}_{n}(\mathbb{C})_{+}:=\{b\otimes c\mid b\in\mathbb{M}_{m}(\mathbb{C})_{+},\,c\in\mathbb{M}_{n}(\mathbb{C})_{+}\}$. Based on the results in this section, we obtain a characterization of separable elements, which is a generalization of the quantum de Finetti theorem.
@@
Let $l$ be a strictly positive integer, $\rho\in\mathbb{M}_{n}(\mathbb{C})^{*}_{+}$ and $a\in(\mathbb{M}_{m}(\mathbb{C})\otimes\mathbb{M}_{n}(\mathbb{C}))_{+}$. An element $b$ in $(\mathbb{M}_{m}(\mathbb{C})\otimes(\mathbb{M}_{n}(\mathbb{C})^{\otimes l})^{S_{l}})_{+}$ a $l$-sub-extension of $a$ (with respect to $\rho$) if
$(\mathrm{id}_{m}\otimes\mathrm{id}_{n}\otimes\rho^{\otimes(l-1)})(b)\leq a.$
If $a$ has a $l$-sub-extension, then $a$ is said to be $l$-sub-extendable (with respect to $\rho$). If the inequalities of this definition are replaced by equalities, the notion of sub-extension (resp. sub-extendable) becomes extension (resp. extendable).
This notion is quite important as a means of characterizing separable states, and detecting entanglement. We refer to [7] for papers laying the foundations of such applications.
@@
    $a$ is $l$-sub-extendable with respect to every faithful linear functional on $\mathbb{M}_{n}(\mathbb{C})$ and for every integer $l\geq 1$.
@@
Let $\rho,\nu\in\mathbb{M}_{n}(\mathbb{C})_{+}^{*}$ be faithful. Then, there exists $r>0$ such that $r\nu\leq\rho$. Thus, if $b$ is a $l$-sub-extension of $a$ with respect to $\rho$, then $r^{l-1}b$ is an $l$-sub-extension of $a$ with respect to $\nu$. Therefore, (2) and (3) are equivalent.
To show that (1) and (2) are equivalent, let us first note that if $a=b\otimes c$ for some $b\in\mathbb{M}_{m}(\mathbb{C})_{+}$ and $c\in\mathbb{M}_{n}(\mathbb{C})_{+}$, then we have $a=(\mathrm{id}_{m}\otimes\mathrm{id}_{n}\otimes\rho^{\otimes(l-1)})(\rho(c)^{-(l-1)}b\otimes c^{\otimes l})$, i.e., $a$ is $l$-sub-extendable for all $l\geq 1$. Hence, if $a$ is separable, it is clearly $l$-sub-extendable for all $l\geq 1$, and consequently (1) implies (2).
Conversely, we assume that there exist $l$-sub-extensions $x_{l}$ of $a$ for all $l\geq 1$. For any $k\geq l$ we define